OpenStudy (anonymous):

What are the equations of motion for an free falling object which falls in spiral fashion? Shall we use same classic mechanics equations?

OpenStudy (abb0t):

Assuming no air resistance, you would use: \[v(t) = -g t + v_0\] \[y(t) = - \frac{ 1 }{ 2 }g t^2 + v_0 t + y_0\] where: v0 = initial velocity (m/s). v(t) = vertical velocity with respect to time (m/s). y0 = initial altitude (m). y(t) = altitude with respect to time (m). t = time (s). g = gravity (9.8 m/s^2)

OpenStudy (anonymous):

how can a free object fall in SPIRAL fashion???

OpenStudy (abb0t):

air resistance?

OpenStudy (anonymous):

what sort of air resistance would give you a perfect spiral motion... and besides whatever it is.. in that case.. the equations you gave will not hold good :P

OpenStudy (abb0t):

i didn't read spiral motion. I just saw free-fall.

OpenStudy (abb0t):

In that case, I don't know an equation for a spiral free-fall. And I don't know anything that would spiral down in a "perfect" spiral motion from free-fall...

OpenStudy (anonymous):

let as assume an object falls due to gravity at the height of 100mm. If we provide some guides like a spring,what can be the acceleration, velocity Shall we apply Newton's equations v=u+at s=ut+1/2at^2 u = initial velocity v= final velocty s= distance travelled